Photochronograph Corporation (PC) manufactures time series photographic equipment. It is currently at its target debt−equity ratio of .80. It’s considering building a new $50 million manufacturing facility. This new plant is expected to generate aftertax cash flows of $6.2 million in perpetuity. The company raises all equity from outside financing. There are three financing options: 1. A new issue of common stock: The flotation costs of the new common stock would be 8 percent of the amount raised. The required return on the company’s new equity is 14 percent. 2. A new issue of 20-year bonds: The flotation costs of the new bonds would be 4 percent of the proceeds. If the company issues these new bonds at an annual coupon rate of 8 percent, they will sell at par. 3. Increased use of accounts payable financing: Because this financing is part of the company’s ongoing daily business, it has no flotation costs, and the company assigns it a cost that is the same as the overall firm WACC. Management has a target ratio of accounts payable to long-term debt of .15. (Assume there is no difference between the pretax and aftertax accounts payable cost.) What is the NPV of the new plant? Assume that PC has a 35 percent tax rate. (Enter your answer in dollars, not millions of dollars, i.e. 1,234,567. Do not round intermediate calculations and round your final answer to the nearest whole dollar amount.)
Answer #1
Photochrono Weight Cost
Equity 55.56% 15.22%
Bonds 38.65% 8.33%
Acct. Payable 5.80% 11.20%
WACC 11.20%
11.20%

Cost of equity after flotation cost, ke = 14% / (1 - 8%) = 15.22%

Cost of debt after flotation cost, kd = 8% / (1 - 4%) = 8.33%

Cost of accounts payable, kap = WACC

Now, as D/E = 0.8, Weight of equity, we = 1 - D / (D + E) = 1 - 0.8 / 1.8 = 55.56%

and weight of total debt = D / (D + E) = 0.8 / 1.8 = 44.44%

Now, Accounts payable to long-term debt = 0.15 => weight of ap, wap = 0.15 / (1 + 0.15) x 44.44% = 5.80%

Weight of long term debt, wd = 44.44% - 5.80% = 38.65%

WACC = we x ke + wap x kap + wd x kd x (1 - tax)

=> WACC = 55.56% x 15.22% + 5.8% x WACC + 38.65% x 8.33% x (1 - 35%)

=> WACC = 11.20%

Now, NPV = - Investment + Cash Flows / WACC

= - 50 + 6.2 / 11.20%

= $5,374,046